OpenStudy (anonymous):

Solve 4x2 = 12x + 40.

OpenStudy (anonymous):

Solve 4x^2 = 12x + 40.

OpenStudy (anonymous):

x² - 3x = 10 x² - 3x - 10 = 0 (x - 5)(x + 2) = 0 x = 5, x = -2

OpenStudy (gorv):

4x2-12x-40=0 take 4 common 4(x2-3x-10)=0 x2-3x-10=0